# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 4.1. Group theoretical Hasse invariants
We come back to the setting of subsection 3.1. Let $(G,\mu)$ be a cocharacter datum over $\mathbb{F}_{p}$, this means that $G$ is a connected reductive group over $\mathbb{F}_{p}$ and $\mu$ a cocharacter of $G$ defined over a finite field $\kappa$ over $\mathbb{F}_{p}$. Recall that $(P_{+},P_{-})$ is the pair of opposite parabolic subgroups of $G_{\kappa}$ determined by $\mu$, and $L:=P_{-}\cap P_{+}$ the Levi subgroup of $G_{\kappa}$. Let $U_{+}$ (resp. $U_{-}$) the unipotent radical of $P_{+}$ (resp. $P_{-}$). The zip group $E=E_{G,\mu}$ is the subgroup of $G_{\kappa}\times G_{\kappa}$
$E=\{(p_{+}=lu_{+},p_{-}=l^{(p)}u_{-})\in P_{+}\times P_{-}^{(p)}\mid l\in L,u_{+}\in U_{+},u_{-}\in U_{-}^{(p)}\}.$
@@
${G}\text{-}\mathrm{Zip}^{\mu}_{\kappa}\cong[E\backslash G_{\kappa}].$
Let $k=\overline{\kappa}$ and ${G}\text{-}\mathrm{Zip}^{\mu}={G}\text{-}\mathrm{Zip}^{\mu}_{k}=[E_{k}\backslash G_{k}]$. Given $\lambda\in X^{*}(L)$, which can be viewed as an element of $X^{*}(E)$ through the projection $E\to L$, one can associate a line bundle $\operatorname{\mathcal{V}}(\lambda)$ over $[E_{k}\backslash G_{k}]$ such that
$H^{0}([E_{k}\backslash G_{k}],\operatorname{\mathcal{V}}(\lambda))=\{f:G_{k}\to\operatorname{\mathbb{A}}^{1}_{k}\mid f(gx)=\lambda(g)f(x)\text{ for all }g\in E_{k},x\in G_{k}\}.$
@@
Recall that a cocharacter datum $(G,\mu)$ is Hodge type if there is a symplectic embedding $G\hookrightarrow\operatorname{GSp}_{2g}$ over $\mathbb{F}_{p}$ such that $\mu$ induces the standard minuscule cocharacter of $\operatorname{GSp}_{2g}$. For such datum, there is a Hodge line bundle $\operatorname{\mathcal{V}}(\eta_{\omega})$ over $[E\backslash G_{\kappa}]$ by pull-back from the given symplectic embedding of $G$ (in general $\operatorname{\mathcal{V}}(\eta_{\omega})$ depends on the symplectic embedding). We have the following theorem.
@@
Moreover, if $(G,\mu)$ is Hodge type ($p=2$ allowed), then Hodge line bundle is a Hasse generator.
@@ 4.2. Length Hasse invariants
Keep notations as above. Let $d=\langle 2\rho,\mu\rangle$, where $\rho$ is the half sum of positive (absolute) roots of $G$. Recall the subsets $G_{w}\subset G$ as in Theorem 3.5 attached to $w\in{}^{J}W$. For any integer $0\leq j\leq d$, define
@@
with the reduced subscheme structure. It is called the $j$-th length stratum of $G$. By [18], we have the following identities:
@@
There exists an integer $N^{\prime}\geq 1$ such that for each $0\leq j\leq d$, a section $h_{j}\in H^{0}([E\backslash\overline{G}_{j}],\operatorname{\mathcal{V}}(\eta_{\omega})^{N^{\prime}})$ such that $D(h_{j})=[E\backslash G_{j}]$.
@@
Now we assume that $S$ is a $k$-scheme and $\zeta:S\rightarrow{G}\text{-}\mathrm{Zip}^{\mu}$ is a morphism of stacks (may not smooth). For a character $\lambda\in X^{\ast}(L)$, write $\mathcal{V}_{S}(\lambda)=\zeta^{\ast}(\mathcal{V}(\lambda))$. For $w\in{}^{J}W$ and $j\in\{0,\ldots,d\}$, we define the locally-closed subsets of $S$: $S_{w},S_{w}^{\ast},S_{j},S_{j}^{\ast}$ as preimage of $[E\backslash G_{w}],[E\backslash\overline{G}_{w}],[E\backslash G_{j}],[E\backslash\overline{G}_{j}]$ respectively, equipped with reduced subscheme structure from $S$.
@@ 4.3. Hasse invariants on splitting models
$\zeta:S\to{\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}}\text{-}\mathrm{Zip}^{\mu}.$
For each $w\in{}^{J}W$, let $S_{w}=\zeta^{-1}(w)$ the EO stratum of $S$. By our construction, $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$ is the similitude group of the lattice $\Lambda^{\mathrm{spl}}_{0}$ with a pairing induced from that on $\Lambda$, hence the pair $(\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}},\mu)$ is of Hodge type and the general theory of Hasse invariants applies.
@@
There exists an integer $N\geq 1$ such that for every $w\in{}^{J}W$, there exists a section $h_{w}\in H^{0}([E\backslash\overline{\operatorname{\mathcal{G}}_{0,w}^{\mathrm{spl}}}],\operatorname{\mathcal{V}}(N\eta_{\omega}))$ whose non-vanishing locus is precisely $[E\backslash\operatorname{\mathcal{G}}_{0,w}^{\mathrm{spl}}]$.
@@ 4.4. Example: Hilbert modular varieties
    ¯$\underline{\operatorname{\mathscr{F}}_{\bullet}}=(\operatorname{\mathscr{F}}_{i,j}^{l})_{1\leq i\leq r,1\leq j\leq f_{i},0\leq l\leq e_{i}}$, each $\operatorname{\mathscr{F}}_{i,j}^{l}$ is a locally free sheaf over $S$ such that
@@
For each $i,j,l$, let $h_{i,j}^{l}$ be the partial Hasse invariant defined by [54], which is a section of certain line bundle over $S$. For a point $x=(\underline{A},\underline{\operatorname{\mathscr{F}}_{\bullet}})\in S(k)$, the construction of partial Hasse invariants shows that
@@
The vanishing loci of these $h_{i,j}^{l}$ cut out the stratification in their paper. We claim that their stratification coincides with the Ekedahl-Oort stratification defined in this paper.
@@
of type $\mu_{i}=(1,0)^{e_{i}f_{i}}$ over $S$. The maps $V_{i,j}^{l}:\operatorname{\mathcal{M}}_{i,j}^{l}\to\operatorname{\mathcal{M}}_{i,j}^{l-1}$ (for $l\geq 2$) and $V_{i,j}^{1}:\operatorname{\mathcal{M}}_{i,j}^{1}\to\operatorname{\mathcal{M}}_{i,j-1}^{e_{i},(p)}$ (for $l=1$) induces maps
@@
by restriction to $\omega$. Such morphisms give sections of $\omega_{i,j}^{l-1}\bigotimes(\omega_{i,j}^{1})^{-1}$ and $\omega_{i,j-1}^{e_{i},\otimes p}\bigotimes(\omega_{i,j}^{1})^{-1}$ respectively, which are the partial Hasse invariants of Reduzzi-Xiao. More precisely, $\omega_{i,j}^{l}$ is just a Frobenius twisted version of $\omega_{\tau_{\mathfrak{p}_{i},j}^{l}}$ in [54], i.e. $(\omega_{i,j}^{l})^{(p)}=\omega_{\tau_{\mathfrak{p}_{i},j}^{l}}$. Now by [25]*Proposition 5.2.3, every codimension one closed Ekedahl-Oort stratum of zip stacks can be cut out from zip partial Hasse invariants. So we at least have codimension one closed EO strata coincide with that in Reduzzi-Xiao’s definition.
In Hilbert case, any EO stratum can be cut out from co-dimension one strata as the associated Weyl group is isomorphic to $(\operatorname{\mathbb{Z}}/2\operatorname{\mathbb{Z}})^{n}$. So all the EO strata of splitting models of Hilbert modular varieties coincide with the strata constructed by the partial Hasse invariants in [54].
@@ 4.5. Example: Hilbert-Siegel case
the similitude subgroup of $\operatorname{Res}_{L|\mathbb{Q}}\operatorname{GSp}_{2g}$. In fact one can also work with Shimura varieties associated to the sightly larger group $\operatorname{Res}_{L|\mathbb{Q}}\operatorname{GSp}_{2g}$, which is of abelian type, by considering some suitable group quotient which does not change the geometry at $p$, see for example [6, 58, 14, 29]. Fix a prime to $p$ level $K^{p}\subset G(\mathbb{A}_{f}^{p})$ which is sufficiently small. Let $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ be the splitting model over $\operatorname{\mathcal{O}}_{F}$ corresponds to $G$. Note that when $g=1$, we return to the Hilbert case in the last subsection.
@@
Recall that the EO index set of $\operatorname{GSp}_{2g}$ is $\{0,1\}^{g}$ (see [62, §5.4]), so the EO index set ${}^{J}W$ of $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}$ can be identified with ($(\{0,1\}^{g})^{n}$, under the isomorphism $\operatorname{\mathcal{G}}_{0,\kappa,der}^{\mathrm{spl}}\otimes_{\operatorname{\mathbb{F}}_{p}}k\simeq(\mathrm{Sp}_{2g})^{n}$. Given $x=(\underline{A},\underline{\operatorname{\mathscr{F}}_{\bullet}})\in\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}(k)$, the EO type of $x$ can be written as a tuple $a=(a_{i,j}^{l})$, where $a_{i,j}^{l}\in\{0,1\}^{g}$. For each index $i,j,l$, by the construction in subsection 3.2 at the point $x$, we have an $F$-zip structure at the standard tuple $(\Lambda_{i,j}^{l}\otimes\kappa,C_{0,i,j}^{l},D_{0,i,j}^{l})$ (of type $\mu_{i,j}^{l}$) by trivializing the filtration
@@
Let $h$ be the product of $h_{i}$, which is a section of the product of the line bundle $\det(\omega_{i}^{e_{i}})^{\otimes(p-1)}$. It coincides with the $\mu$-ordinary Hasse invariant coming from the zip stack.
@@ 4.6. Example: unitary Shimura varieties
[…]
@@
On the algebraically closed field $k$, the group $\operatorname{\mathcal{G}}_{0,k}^{\mathrm{spl}}$ is just the associated similitude subgroup of
@@
where $S_{d_{i}}$ is the permutation group of $d_{i}$ elements, and $(S_{d_{i,j}^{l}}\times S_{d_{i}-d_{i,j}^{l}})$ is the subgroup of $S_{d_{i}}$ consists of the permutation of first $d_{i,j}^{l}$ element times and the permutation of last $d_{i}-d_{i,j}^{l}$ element. For a precise description of the minimal length representatives for the left coset $(S_{d_{i,j}^{l}}\times S_{d_{i}-d_{i,j}^{l}})\setminus S_{d_{i}}$, see [41] subsection 2.6. The partial order $\preceq$ on ${}^{J}W$ is given by the product of the partial orders on $W_{i}^{l}$. We remark that different type of $i$ induces different Frobenius actions on $W_{i}^{l}$, which induces different partial orders $\preceq$ on $W_{i}^{l}$. The construction of Hasse invariants is the same as above.
@@ 5.1. Modifications of moduli spaces and integral models
thus composed with the closed immersion $\operatorname{\mathscr{A}}_{K^{p}}\subset\operatorname{\mathscr{A}}^{\operatorname{naive}}_{K^{p}}$ we get also a closed immersion $\mathsf{M}_{K}\subset\operatorname{\mathscr{A}}^{\operatorname{naive}}_{K^{p}}$. On generic fibers, we have an open and closed embedding of $E$-schemes (cf. [31] Lemma 1.4.4.2)
@@
Let (¯,¯$(\underline{\mathcal{H}},\underline{\operatorname{\mathscr{F}}},\underline{\iota})$ be the polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{\mathsf{M}_{K}}$-modules (see Definition A.3) associated to the pullback of the universal object over $\operatorname{\mathscr{A}}^{\operatorname{naive}}$. Here in fact $\operatorname{\mathscr{F}}=\omega$. By the notation of [35] Proposition 2.3.10, we define the corresponding splitting model as the moduli scheme of splitting structures of (¯,¯$(\underline{\mathcal{H}},\underline{\operatorname{\mathscr{F}}},\underline{\iota})$ over $\mathsf{M}_{K}\otimes\operatorname{\mathcal{O}}_{F}$
@@
The difference of these schemes is bounded by the size of failure of Hasse principle (cf. [31] Remark 1.4.3.12). All the results in previous sections still hold for $\mathsf{M}_{K}^{\mathrm{spl}}$ and $\mathsf{M}_{K}$, as in fact our previous constructions can be done for the related integral models of the associated Shimura varieties.

### v3

@@ 4.1. Group theoretical Hasse invariants
We come back to the setting of subsection 3.1. Let $(G,\mu)$ be a cocharacter datum over $\mathbb{F}_{p}$, this means that $G$ is a connected reductive group over $\mathbb{F}_{p}$ and $\mu$ is a cocharacter of $G$ defined over a finite field $\kappa$ over $\mathbb{F}_{p}$. Recall that $(P_{+},P_{-})$ is the pair of opposite parabolic subgroups of $G_{\kappa}$ determined by $\mu$, and $L:=P_{-}\cap P_{+}$ the Levi subgroup of $G_{\kappa}$. Let $U_{+}$ (resp. $U_{-}$) the unipotent radical of $P_{+}$ (resp. $P_{-}$). The zip group $E_{G,\mu}$ is the subgroup of $G_{\kappa}\times G_{\kappa}$
$E_{G,\mu}=\{(p_{+}=lu_{+},p_{-}=l^{(p)}u_{-})\in P_{+}\times P_{-}^{(p)}\mid l\in L,u_{+}\in U_{+},u_{-}\in U_{-}^{(p)}\}.$
@@
${G}\text{-}\mathrm{Zip}^{\mu}_{\kappa}\cong[E_{G,\mu}\backslash G_{\kappa}].$
[…]
$H^{0}([E_{G,\mu,k}\backslash G_{k}],\operatorname{\mathcal{V}}(\lambda))=\{f:G_{k}\to\operatorname{\mathbb{A}}^{1}_{k}\mid f(gx)=\lambda(g)f(x)\text{ for all }g\in E_{G,\mu,k},x\in G_{k}\}.$
@@
Recall that a cocharacter datum $(G,\mu)$ is Hodge type if there is a symplectic embedding $G\hookrightarrow\operatorname{GSp}_{2g}$ over $\mathbb{F}_{p}$ such that $\mu$ induces the standard minuscule cocharacter of $\operatorname{GSp}_{2g}$. For such datum, there is a Hodge line bundle $\operatorname{\mathcal{V}}(\eta_{\omega})$ over $E_{G,\mu}\backslash G_{\kappa}]$ by pull-back from the given symplectic embedding of $G$ (in general $\operatorname{\mathcal{V}}(\eta_{\omega})$ depends on the symplectic embedding). We have the following theorem.
@@
Moreover, if $(G,\mu)$ is Hodge type ($p=2$ allowed), then the Hodge line bundle is associated to a Hasse generator.
@@ 4.2. Length Hasse invariants
Keep notations as above. Let $d=\langle 2\rho,\mu\rangle$, where $\rho$ is the half sum of positive (absolute) roots of $G$. Recall the subsets $G_{w}\subset G_{k}$ as in Theorem 3.5 attached to $w\in{}^{J}W$. For any integer $0\leq j\leq d$, define
@@
with the reduced subscheme structure. It is called the $j$-th length stratum of $G_{k}$. By [20], we have the following identities:
@@
[…]
@@
[…]
@@ 4.3. Hasse invariants on splitting models
$\zeta:S\to{\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}}\text{-}\mathrm{Zip}^{\mu}_{k}.$
For each $w\in{}^{J}W$, let $S_{w}=\zeta^{-1}(w)$ be the EO stratum of $S$. By our construction, $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$ is the similitude group of the lattice $\Lambda^{\mathrm{spl}}_{0}$ with a pairing induced from that on $\Lambda$, hence the pair $(\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}},\mu)$ is of Hodge type and the general theory of Hasse invariants applies. For simplicity, here in the following we denote $\mathsf{E}=E_{\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}},\mu,k}$.
@@
There exists an integer $N\geq 1$ such that for every $w\in{}^{J}W$, there exists a section $h_{w}\in H^{0}([\mathsf{E}\backslash\overline{\operatorname{\mathcal{G}}_{0,w}^{\mathrm{spl}}}],\operatorname{\mathcal{V}}(N\eta_{\omega}))$ whose non-vanishing locus is precisely $\mathsf{E}\backslash\operatorname{\mathcal{G}}_{0,w}^{\mathrm{spl}}]$.
@@ 4.4. Example: Hilbert modular varieties
    $\underline{\operatorname{\mathscr{F}}_{\bullet}}=(\operatorname{\mathscr{F}}_{i,j}^{l})_{1\leq i\leq r,1\leq j\leq f_{i},0\leq l\leq e_{i}}$, each $\operatorname{\mathscr{F}}_{i,j}^{l}$ is a locally free sheaf over $S$ such that
@@
For each $i,j,l$, let $h_{i,j}^{l}$ be the partial Hasse invariant defined by [56], which is a section of a certain line bundle over $S$. For a point $x=(\underline{A},\underline{\operatorname{\mathscr{F}}_{\bullet}})\in S(k)$, the construction of partial Hasse invariants shows that
@@
The vanishing loci of these $h_{i,j}^{l}$ cut out the stratification in [56]. We claim that their stratification coincides with the Ekedahl-Oort stratification defined in this paper.
@@
of type $\mu_{i}=(1,0)^{e_{i}f_{i}}$ over $S$. The maps $V_{i,j}^{l}:\operatorname{\mathcal{M}}_{i,j}^{l}\to\operatorname{\mathcal{M}}_{i,j}^{l-1}$ (for $l\geq 2$) and $V_{i,j}^{1}:\operatorname{\mathcal{M}}_{i,j}^{1}\to\operatorname{\mathcal{M}}_{i,j-1}^{e_{i},(p)}$ (for $l=1$) induce maps
@@
by restriction to $\omega$. Such morphisms give sections of $\omega_{i,j}^{l-1}\bigotimes(\omega_{i,j}^{1})^{-1}$ and $\omega_{i,j-1}^{e_{i},\otimes p}\bigotimes(\omega_{i,j}^{1})^{-1}$ respectively, which are the partial Hasse invariants of Reduzzi-Xiao. More precisely, $\omega_{i,j}^{l}$ is just a Frobenius twisted version of $\omega_{\tau_{\mathfrak{p}_{i},j}^{l}}$ in [56], i.e. $(\omega_{i,j}^{l})^{(p)}=\omega_{\tau_{\mathfrak{p}_{i},j}^{l}}$. Now by [27]*Proposition 5.2.3, every codimension one closed Ekedahl-Oort stratum of zip stacks can be cut out from zip partial Hasse invariants. So we at least have codimension one closed EO strata coincide with those in Reduzzi-Xiao’s definition.
In the Hilbert case, any EO stratum can be cut out from codimension one strata as the associated Weyl group is isomorphic to $(\operatorname{\mathbb{Z}}/2\operatorname{\mathbb{Z}})^{n}$. So all the EO strata of splitting models of Hilbert modular varieties coincide with the strata constructed by the partial Hasse invariants in [56].
@@ 4.5. Example: Hilbert-Siegel case
the similitude subgroup of $\operatorname{Res}_{L|\mathbb{Q}}\operatorname{GSp}_{2g}$. In fact one can also work with Shimura varieties associated to the sightly larger group $\operatorname{Res}_{L|\mathbb{Q}}\operatorname{GSp}_{2g}$, which is of abelian type, by taking some suitable group quotient which does not change the geometry at $p$, see for example [8, 61, 16, 31]. Fix a prime to $p$ level $K^{p}\subset G(\mathbb{A}_{f}^{p})$ which is sufficiently small. Let $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ be the splitting model over $\operatorname{\mathcal{O}}_{F}$ corresponding to $G$. Note that when $g=1$, we return to the Hilbert case in the last subsection.
@@
Recall that the EO index set of $\operatorname{GSp}_{2g}$ is $\{0,1\}^{g}$ (see [65, §5.4]), so the EO index set $^{J}W$ of $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}$ can be identified with $(\{0,1\}^{g})^{n}$, under the isomorphism $\operatorname{\mathcal{G}}_{0,\kappa,der}^{\mathrm{spl}}\otimes_{\operatorname{\mathbb{F}}_{p}}k\simeq(\mathrm{Sp}_{2g})^{n}$. Given $x=(\underline{A},\underline{\operatorname{\mathscr{F}}_{\bullet}})\in\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}(k)$, the EO type of $x$ can be written as a tuple $a=(a_{i,j}^{l})$, where $a_{i,j}^{l}\in\{0,1\}^{g}$. For each index $i,j,l$, by the construction in subsection 3.2 at the point $x$, we have an $F$-zip structure at the standard tuple $(\Lambda_{i,j}^{l}\otimes\kappa,C_{0,i,j}^{l},D_{0,i,j}^{l})$ (of type $\mu_{i,j}^{l}$) by trivializing the filtrations
@@
Let $h$ be the product of the $h_{i}$, which is a section of the product of the line bundles $\det(\omega_{i}^{e_{i}})^{\otimes(p-1)}$. It coincides with the $\mu$-ordinary Hasse invariant coming from the zip stack.
@@ 4.6. Example: unitary Shimura varieties
Let $V$ be an $L$-vector space of dimension $n$ together with a hermitian form $\langle\cdot,\cdot\rangle$. We assume that this form is not totally definite. Let $G=GU(V,\langle\cdot,\cdot\rangle)$ be the associated reductive group of unitary similitudes over $\mathbb{Q}$. Let $(p_{\tau},q_{\tau})_{\tau\in I}$ be the signature of $G_{\mathbb{R}}$ so that $G_{\mathbb{R}}=G(\prod_{\tau\in I}U(p_{\tau},q_{\tau}))$ and we get a standard $h:\operatorname{Res}_{\mathbb{C}|\mathbb{R}}\mathbb{G}_{m}\rightarrow G_{\mathbb{R}}$. If we write $G_{1}=U(V,\langle\cdot,\cdot\rangle)$ as the corresponding unitary group over $\mathbb{Q}$, then $G_{1}=\operatorname{Res}_{L^{+}|\mathbb{Q}}U$, where $U$ is the unitary group over $L^{+}$ defined by $(V,\langle\cdot,\cdot\rangle)$.
@@
Over the algebraically closed field $k$, the group $\operatorname{\mathcal{G}}_{0,k}^{\mathrm{spl}}$ is just the associated similitude subgroup of
@@
where $S_{d_{i}}$ is the permutation group of $d_{i}$ elements, and $(S_{d_{i,j}^{l}}\times S_{d_{i}-d_{i,j}^{l}})$ is the subgroup of $S_{d_{i}}$ consisting of the permutations of the first $d_{i,j}^{l}$ elements and the permutations of the last $d_{i}-d_{i,j}^{l}$ elements. For a precise description of the minimal length representatives for the left coset $(S_{d_{i,j}^{l}}\times S_{d_{i}-d_{i,j}^{l}})\setminus S_{d_{i}}$, see [43] subsection 2.6. The partial order $\preceq$ on $^{J}W$ is given by the product of the partial orders on $W_{i}^{l}$. We remark that different types of $i$ induce different Frobenius actions on $W_{i}^{l}$, which induces different partial orders $\preceq$ on $W_{i}^{l}$. The construction of Hasse invariants is the same as above.
@@ 5.1. Modifications of moduli spaces and integral models
thus composing with the closed immersion $\operatorname{\mathscr{A}}_{K^{p}}\subset\operatorname{\mathscr{A}}^{\operatorname{naive}}_{K^{p}}$ we get also a closed immersion $\mathsf{M}_{K}\subset\operatorname{\mathscr{A}}^{\operatorname{naive}}_{K^{p}}$. On generic fibers, we have an open and closed embedding of $E$-schemes (cf. [33] Lemma 1.4.4.2)
@@
Let $(\underline{\mathcal{H}},\underline{\operatorname{\mathscr{F}}},\underline{\iota})$ be the polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{\mathsf{M}_{K}}$-modules (see Definition A.3) associated to the pullback of the universal object over $\operatorname{\mathscr{A}}^{\operatorname{naive}}$. Here in fact $\operatorname{\mathscr{F}}=\omega$. By the notation of [37] Proposition 2.3.10, we define the corresponding splitting model as the moduli scheme of splitting structures of $(\underline{\mathcal{H}},\underline{\operatorname{\mathscr{F}}},\underline{\iota})$ over $\mathsf{M}_{K}\otimes\operatorname{\mathcal{O}}_{F}$
@@
The difference of these schemes is bounded by the size of failure of the Hasse principle (cf. [33] Remark 1.4.3.12). All the results in previous sections still hold for $\mathsf{M}_{K}^{\mathrm{spl}}$ and $\mathsf{M}_{K}$, as in fact our previous constructions can be done for the related integral models of the associated Shimura varieties.